OpenStudy (anonymous):

Please need help with this problem..The difference between 9 times a number and 16 is 74. find the number.

OpenStudy (ajprincess):

Let the number be x. Then 9x-16=74 Can u find x @lopez23?

OpenStudy (anonymous):

9x-16=74 9x=90 x=10

OpenStudy (anonymous):

thanks alot guys.

OpenStudy (ajprincess):

welcome:)

OpenStudy (anonymous):

let x be the number 9x-16=74 9x=90 x=10

OpenStudy (anonymous):

you're welcome